# Spin and energy diffusion vs. subdiffusion in disordered spin chains

### v1

@@ VI Summary
Besides the more frequently discussed spin diffusion, we investigate in this paper the energy diffusion within the XXZ Heisenberg spin chain, which is subject to random local fields. The memory-function approach in the hydrodynamic regime allows the derivation of appropriate Einstein relations, expressing diffusion response in terms of spin and energy current correlations, respectively. Following numerical results on finite systems with up to $L=28$ sites for both dynamical diffusivities ${\cal D}_{s,e}(\omega)$ as well for d.c. diffusion constants ${\cal D}^{0}_{s,e}$ reveal some essential similarities but as well differences. Static ${\cal D}^{0}_{s,e}$ show strong exponential-like dependence on the field strength $W$ but also pronounced dependence and variation with the actual random field configurations $w_{l}$. While energy diffusion is evidently faster, i.e. ${\cal D}^{0}_{e}>{\cal D}^{0}_{s}$ at weak $W<1.5$ [49], both diffusion become (up to numerical resolution) nearly equal at larger $W>2.0$, both governed by the incoherent spin transport. This conclusion is important in relation to frequently discussed Thouless time determined by the slowest diffusion $\tau_{\mathrm{Th}}=L^{2}/{\cal D}^{0}$, where our results indicate that at large disorders, both diffusions are equally important.
@@
Due to modest system sizes $L\leq 28$, the values of the d.c. diffusion constants ${\cal D}^{0}_{s,e}$ show large fluctuations between different field configurations at the same effective strength $W$. We can partly reduce these variations (at the same $W$) by fixing/renormalizing the random amplitudes $\sum_{l}\tilde{\eta}^{2}_{l}=1/3$. Still, at larger $W$, the distributions of ${\cal D}^{0}_{s,e}$ are becoming non-Gaussian, i.e., closer to log-normal distribution, even with pronounced tails.

### v2

@@ VI Summary
Besides the more frequently discussed spin diffusion, we investigate in this paper the energy diffusion within the XXZ Heisenberg spin chain, which is subject to random local fields. The memory-function approach in the hydrodynamic regime allows the derivation of appropriate Einstein relations, expressing diffusion response in terms of spin and energy current correlations, respectively. Following numerical results on finite systems with up to $L=28$ sites for both dynamical diffusivities ${\cal D}_{s,e}(\omega)$ as well as for d.c. diffusion constants ${\cal D}^{0}_{s,e}$ reveal some essential similarities but as well differences. Static ${\cal D}^{0}_{s,e}$ shows strong exponential-like dependence on the field strength $W$ but also pronounced dependence and variation with the actual random field configurations $w_{l}$. While energy diffusion is evidently faster, i.e. ${\cal D}^{0}_{e}>{\cal D}^{0}_{s}$ at weak $W<1.5$ [48], both diffusion become (up to numerical resolution) nearly equal at larger $W>2.0$, both governed by the incoherent spin transport. This conclusion is important in relation to the frequently discussed Thouless time determined by the slowest diffusion $\tau_{\mathrm{Th}}=L^{2}/{\cal D}^{0}$, where our results indicate that at large disorders, both diffusions are equally important.
@@
Due to modest system sizes $L\leq 28$, the values of the d.c. diffusion constants ${\cal D}^{0}_{s,e}$ show large fluctuations between different field configurations at the same effective strength $W$. We can partly reduce these variations (at the same $W$) by fixing/renormalizing the random amplitudes $\sum_{l}\tilde{\eta}^{2}_{l}=1/3$. Still, at larger $W$, the distributions of ${\cal D}^{0}_{s,e}$ are becoming non-Gaussian, i.e., closer to a log-normal distribution, even with pronounced tails.
@@
The data that support the findings of this article are openly available [67].